OpenStudy (anonymous):

Evaluate arccsc√2. A. {π/3± 2πn, 5π/3± 2πn} B. {π/4±2nπ, 3π/4± 2nπ} C. {π/6, 5π/6} D. π/3

OpenStudy (anonymous):

once again these are single valued functions so only answer that make senses is the last one

OpenStudy (anonymous):

\[cosec^{-1}\sqrt{2} = y\] \[\sqrt2 = cosecy\] \[y = \frac{\pi}{4}\]

OpenStudy (anonymous):

hope this is clear. you cannot say evaluate \[f(\sqrt{2})\] and then give a set of numbers. must be one number or it isn't a function

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

\[\sec(\frac{\pi}{3})=\frac{1}{\sin(\frac{\pi}{3})}=\frac{2}{\sqrt{3}}\]

OpenStudy (anonymous):

so that answer is not right. you have no correct answer here. where did the question come from?

OpenStudy (anonymous):

online quiz

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

then put in B and call tech support and tell them the question is in error.

OpenStudy (anonymous):

god no wonder students are so confused.

OpenStudy (anonymous):

haha, alright

OpenStudy (anonymous):

sorry